# Provable test-time adaptivity and distributional robustness of in-context learning

### v1

@@ Appendix B Proof of Proposition 1
Hence,
$\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})$ $-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\hskip-7.00032pt\overset{(i)}{=}\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\hskip-8.72289pt\overset{(ii)}{\leq}4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr].$
Here, $(ii)$ follows from (15), and $(i)$ follows since
@@
$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X)\bigr\},$
@@
$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\bigr\}=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\,|\,X\bigr\}\bigr]=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}.$
This proves the claim. ∎
@@ C.1 Proof of Theorem 2
$Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\coloneqq\begin{pmatrix}X_{1}^{\top}&Y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ X_{n}^{\top}&Y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ X^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$ (17)
@@
$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,X_{1},Y_{1},\ldots,X_{N},Y_{N},X,Y\biggr)\biggr\}\leq\delta.$
@@
$\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\biggr)$
$\hskip 113.81102pt=\mathbb{E}\biggl\{\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\,\biggm|\,X_{1},\ldots,X_{n}\biggr)\biggr\}=0.$
@@
$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}$
$\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}+\frac{\delta}{3}\leq\delta.$
@@
Define a Borel measurable function $f:\mathbb{R}^{d_{\mathrm{model}}}\to\mathbb{R}^{d_{\mathrm{model}}}$ by $f(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)\coloneqq f_{1}(x,y)-(x,y,\bm{0}_{d_{\mathrm{model}}-d-2},0)$ and $f(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)\coloneqq f_{2}(x)-(x,0,\bm{0}_{d_{\mathrm{model}}-d-2},1)$ for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ and set $f$ to be zero otherwise. Let $\mu_{1}$ denote the distribution of $(X,Y)$ and let $\mu_{2}$ denote the marginal distribution of $X$. Define a distribution $\mu$ on $\mathbb{R}^{d_{\mathrm{model}}}$ by
$\mu\biggl(\prod_{j=1}^{d_{\mathrm{model}}}A_{j}\biggr)\coloneqq\frac{\mu_{1}(\prod_{j=1}^{d+1}A_{j})\mathbbm{1}_{\{0\in A_{d_{\mathrm{model}}}\}}+\mu_{2}(\prod_{j=1}^{d}A_{j})\mathbbm{1}_{\{1\in A_{d_{\mathrm{model}}}\}}}{2}\cdot\prod_{j=d+2}^{{d_{\mathrm{model}}}-1}\mathbbm{1}_{\{0\in A_{j}\}},$
@@
$\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ $(X,Y,\bm{0}_{{d_{\mathrm{model}}}-d-2},0)-f_{1}(X,Y)\bigr\|_{\infty}>\epsilon\Bigr)$
$+\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(X,0,\bm{0}_{{d_{\mathrm{model}}}-d-2},1)-f_{2}(X)\bigr\|_{\infty}>\epsilon\Bigr)\leq\frac{\delta}{N+1}.$
@@
$\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\bigr]_{i,{d_{\mathrm{model}}}}=\bigl[Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})}\bigr]_{i,{d_{\mathrm{model}}}}=\mathbbm{1}_{\{i=n+1\}}=\bigl[U^{(n)}\bigr]_{i,{d_{\mathrm{model}}}}\quad\text{for }i\in[n+1].$ (18)
Next, we define $Q^{(2)}=K^{(2)}\coloneqq\bm{0}_{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ and define $V^{(2)}\in\mathbb{R}^{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ by $V^{(2)}_{j,2M+j}\coloneqq 1$ for $j\in[2M]$, $V^{(2)}_{4M+1,4M+1}\coloneqq 1$ and zero otherwise. Let $\theta^{(2)}_{\mathrm{attn}}\coloneqq(Q^{(2)},K^{(2)},V^{(2)})$, $\widetilde{Z}^{(n)}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ and $\bar{Z}^{(n)}\coloneqq\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}(\widetilde{Z}^{(n)})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$. Then by construction and (18), the $(n+1)$th row of $\bar{Z}^{(n)}$ is given by
@@ D.1 Function spaces
$\Delta_{h}^{r}(f)(x)=\mathbbm{1}_{\{x+rh\in[0,1]^{d}\}}\sum_{k=0}^{r}\binom{r}{k}(-1)^{r-k}f(x+kh).$
Note that $\Delta_{h}^{1}(f)(x)=\{f(x+h)-f(x)\}\mathbbm{1}_{\{x+h\in[0,1]^{d}\}}$ and $\Delta_{h}^{r}(f)(x)=\Delta_{h}^{1}\bigl(\Delta_{h}^{r-1}(f)(x)\bigr)$ if $x+rh\in[0,1]^{d}$. Now, for $p\in[1,\infty]$, $f\in L^{p}([0,1]^{d})$ and $t\geq 0$, define the $r$th modulus of smoothness
@@ D.2 Wavelets
Further, for $\tau=(\tau_{1},\ldots,\tau_{d})^{\top}\in\{0,1\}^{d}\setminus\{\bm{0}_{d}\}$, $\ell\in[{\ell_{0}}:\infty)$, $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
@@
$K\coloneqq[0:2^{\ell_{0}}-1]^{d}\quad\text{and}\quad\Gamma_{\ell}\coloneqq[0:2^{\ell}-1]^{d}\times\bigl(\{0,1\}^{d}\setminus\{\bm{0}_{d}\}\bigr)\text{ for }\ell\in[\ell_{0}:\infty),$ (30)
@@
$\sum_{k\in K}\mathbbm{1}_{\{\Phi_{k}(x)\neq 0\}}\vee\sup_{\ell\in[\ell_{0}:\infty)}\sum_{\gamma\in\Gamma_{\ell}}\mathbbm{1}_{\{\Psi_{\ell,\gamma}(x)\neq 0\}}\leq C^{\prime}\quad\text{for all }x\in[0,1]^{d}.$ (33)
@@ E.2 Bayesian nonparametric regression
$\hskip-7.00032pt\overset{(i)}{\leq}\mathbb{E}_{P_{0}}\biggl[\int_{\mathcal{P}}\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-8.72289pt\overset{(ii)}{=}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl(\int_{\mathcal{P}}\mathbb{E}_{(X,Y)\sim P_{0}}\bigl[\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\bigr]\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr)$
$\hskip-10.44545pt\overset{(iii)}{\leq}C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl[\int_{\mathcal{P}}d_{\mathrm{H}}^{2}(P,P_{0})\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-9.60333pt\overset{(iv)}{\leq}C_{1}^{\prime}M^{2}\epsilon_{n,\beta}^{2}+2C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,\mathcal{D}_{n}\bigr)\bigr]$
$\hskip-7.88077pt\overset{(v)}{\leq}C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
@@ F.2 Proof of Proposition 6
$g^{\circ}(x)=h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr),$ (55)
@@
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\beta,d,r,c_{0},C_{0}$ and $S$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may in addition depend on $\sigma$. Let $\widetilde{g}_{\beta,r}(x)=\widetilde{g}_{\beta}^{(r)}\bigl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\bigr)$ be defined as in (11) so that $\widetilde{g}_{\beta,r}\sim\widetilde{\pi}_{\beta,r}$. Observe that $B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})\subseteq B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$ (e.g. Giné and Nickl, 2021, Proposition 4.3.10(ii)). When $\beta<1$, the space $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Hölder space $H^{\beta}$ with equivalent norms (e.g. Giné and Nickl, 2021, Proposition 4.3.23), so for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{1}\|x-y\|_{2}^{\beta}$; when $\beta\geq 1$, $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Zygmund space of order 11 with equivalent norms (Triebel, 1983, p. 113), so by Anderson and Pitt (1989, Proposition 2.4) (see also Zygmund, 2002, Chapter II, Theorem 3.4), for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{2}\|x-y\|_{2}\bigl\{\log(1/\|x-y\|_{2})\vee 1\bigr\}$. Therefore, for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$,
@@
$\leq\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}\biggr)-\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|+\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\biggr)\biggr|$
@@
where the second inequality follows from (59), since $\widetilde{g}_{\beta}^{(r)}\in\mathcal{G}_{\beta}^{(r)}\subseteq B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$ and $\bigl\|\frac{(U^{(r)})^{\top}x+\bm{1}_{r}}{2}-\frac{(U^{\circ})^{\top}x+\bm{1}_{r}}{2}\bigr\|_{2}\leq\frac{1}{2}\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}$ on $E$. Now by Lemma 26, we have

### v2

@@ Appendix B Proof of Proposition 1
Further,
@@
$=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})\cdot\mathbb{E}_{P}(Y\,|\,X)\bigr\},$ (15)
@@
$\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\bigr\}=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\mathbb{E}_{P}\bigl\{Y\mathbb{E}_{P}(Y\,|\,X)\,|\,X\bigr\}\bigr]=\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl\{\bigl(\mathbb{E}_{P}(Y\,|\,X)\bigr)^{2}\bigr\}.$ (16)
Hence, we deduce that
$\mathcal{R}_{\mu}^{\mathrm{ICL}}(\widetilde{f}_{T})$ $=\mathbb{E}\mathcal{R}_{\mu}(\widetilde{f}_{T})-\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$
$\hskip-5.92851pt\overset{(i)}{=}\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ (17)
$\hskip-7.34315pt\overset{(ii)}{\leq}4R\sqrt{\bigl\{\chi^{2}(\mu,\pi)+1\bigr\}\mathcal{E}(\widetilde{f}_{T})}+2\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$
where $(i)$ follows from (15) and (16), and $(ii)$ follows from (14). ∎
@@ C.1 Proof of Theorem 2
$Z_{\mathrm{in}}^{(n,d_{\mathrm{model}})}\coloneqq\begin{pmatrix}X_{1}^{\top}&Y_{1}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ X_{n}^{\top}&Y_{n}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ X^{\top}&0&0_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}.$ (19)
@@
$=\mathbb{E}\biggl\{\mathbb{P}\biggl(\bigcup_{n=1}^{N}(A_{n}^{\mathrm{c}}\cup B_{n}^{\mathrm{c}})\,\bigg|\,X_{1},Y_{1},\ldots,X_{N},Y_{N},X,Y\biggr)\biggr\}\leq\delta.$
@@
$\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}\bar{f}_{\xi}\bigl(y_{i},g(x_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\biggr)=\mathbb{E}\biggl\{\mathbb{P}\biggl(\int_{\mathcal{G}}\prod_{i=1}^{n}f_{\xi}\bigl(y_{i},g(X_{i})\bigr)\,\mathrm{d}\widetilde{\pi}(g)=0\,\biggm|\,X_{1},\ldots,X_{n}\biggr)\biggr\}=0.$
@@
$\hskip 113.81102pt\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|f(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}$
$\hskip 113.81102pt\leq\mu\Bigl(\bigl\{x\in K_{0}:\bigl|h(x)-W_{2}^{\top}\rho(W_{1}^{\top}x+v)\bigr|\geq\epsilon\bigr\}\Bigr)+\frac{\delta}{3}+\frac{\delta}{3}\leq\delta.$
@@
Define a Borel measurable function $f:\mathbb{R}^{d_{\mathrm{model}}}\to\mathbb{R}^{d_{\mathrm{model}}}$ by $f(x,y,0_{d_{\mathrm{model}}-d-2},0)\coloneqq f_{1}(x,y)-(x,y,0_{d_{\mathrm{model}}-d-2},0)$ and $f(x,0,0_{d_{\mathrm{model}}-d-2},1)\coloneqq f_{2}(x)-(x,0,0_{d_{\mathrm{model}}-d-2},1)$ for $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ and set $f$ to be zero otherwise. Let $\mu_{1}$ denote the distribution of $(X,Y)$ and let $\mu_{2}$ denote the marginal distribution of $X$. Define a distribution $\mu$ on $\mathbb{R}^{d_{\mathrm{model}}}$ by
$\mu\biggl(\prod_{j=1}^{d_{\mathrm{model}}}A_{j}\biggr)\coloneqq\frac{\mu_{1}(\prod_{j=1}^{d+1}A_{j})\mathbbold{1}_{\{0\in A_{d_{\mathrm{model}}}\}}+\mu_{2}(\prod_{j=1}^{d}A_{j})\mathbbold{1}_{\{1\in A_{d_{\mathrm{model}}}\}}}{2}\cdot\prod_{j=d+2}^{{d_{\mathrm{model}}}-1}\mathbbold{1}_{\{0\in A_{j}\}},$
@@
$\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}$ $(X,Y,0_{{d_{\mathrm{model}}}-d-2},0)-f_{1}(X,Y)\bigr\|_{\infty}>\epsilon\Bigr)$
$+\mathbb{P}\Bigl(\bigl\|\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(X,0,0_{{d_{\mathrm{model}}}-d-2},1)-f_{2}(X)\bigr\|_{\infty}>\epsilon\Bigr)\leq\frac{\delta}{N+1}.$
@@
$\bigl[\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\bigr]_{i,{d_{\mathrm{model}}}}=\bigl[Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})}\bigr]_{i,{d_{\mathrm{model}}}}=\mathbbold{1}_{\{i=n+1\}}=\bigl[U^{(n)}\bigr]_{i,{d_{\mathrm{model}}}}\quad\text{for }i\in[n+1].$ (20)
Next, we define $Q^{(2)}=K^{(2)}\coloneqq 0_{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ and define $V^{(2)}\in\mathbb{R}^{{d_{\mathrm{model}}}\times{d_{\mathrm{model}}}}$ by $V^{(2)}_{j,2M+j}\coloneqq 1$ for $j\in[2M]$, $V^{(2)}_{4M+1,4M+1}\coloneqq 1$ and zero otherwise. Let $\theta^{(2)}_{\mathrm{attn}}\coloneqq(Q^{(2)},K^{(2)},V^{(2)})$, $\widetilde{Z}^{(n)}\coloneqq\mathsf{FFN}_{\theta_{\mathrm{ffn}}^{(1)}}(Z_{\mathrm{in}}^{(n,{d_{\mathrm{model}}})})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$ and $\bar{Z}^{(n)}\coloneqq\mathsf{Attn}_{\theta^{(2)}_{\mathrm{attn}}}(\widetilde{Z}^{(n)})\in\mathbb{R}^{(n+1)\times{d_{\mathrm{model}}}}$. Then by construction and (20), the $(n+1)$th row of $\bar{Z}^{(n)}$ is given by
@@ D.1 Function Spaces
$\Delta_{h}^{r}(f)(x)=\mathbbold{1}_{\{x+rh\in[0,1]^{d}\}}\sum_{k=0}^{r}\binom{r}{k}(-1)^{r-k}f(x+kh).$
Note that $\Delta_{h}^{1}(f)(x)=\{f(x+h)-f(x)\}\mathbbold{1}_{\{x+h\in[0,1]^{d}\}}$ and $\Delta_{h}^{r}(f)(x)=\Delta_{h}^{1}\bigl(\Delta_{h}^{r-1}(f)(x)\bigr)$ if $x+rh\in[0,1]^{d}$. Now, for $p\in[1,\infty]$, $f\in L^{p}([0,1]^{d})$ and $t\geq 0$, define the $r$th modulus of smoothness
@@ D.2 Wavelets
Further, for $\tau=(\tau_{1},\ldots,\tau_{d})^{\top}\in\{0,1\}^{d}\setminus\{0_{d}\}$, $\ell\in[{\ell_{0}}:\infty)$, $k=(k_{1},\ldots,k_{d})^{\top}\in[0:2^{\ell}-1]^{d}$ and $x=(x_{1},\ldots,x_{d})^{\top}\in[0,1]^{d}$, define
@@
$K\coloneqq[0:2^{\ell_{0}}-1]^{d}\quad\text{and}\quad\Gamma_{\ell}\coloneqq[0:2^{\ell}-1]^{d}\times\bigl(\{0,1\}^{d}\setminus\{0_{d}\}\bigr)\text{ for }\ell\in[\ell_{0}:\infty),$ (32)
@@
[…]
@@ E.2 Bayesian Nonparametric Regression
$\hskip-5.92851pt\overset{(i)}{\leq}\mathbb{E}_{P_{0}}\biggl[\int_{\mathcal{P}}\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-7.34315pt\overset{(ii)}{=}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl(\int_{\mathcal{P}}\mathbb{E}_{(X,Y)\sim P_{0}}\bigl[\bigl\{g_{P}(X)-g^{0}(X)\bigr\}^{2}\bigr]\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr)$
$\hskip-8.7578pt\overset{(iii)}{\leq}C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\biggl[\int_{\mathcal{P}}d_{\mathrm{H}}^{2}(P,P_{0})\,\mathrm{d}\pi(P\,|\,\mathcal{D}_{n})\biggr]$
$\hskip-8.05411pt\overset{(iv)}{\leq}C_{1}^{\prime}M^{2}\epsilon_{n,\beta}^{2}+2C_{1}^{\prime}\mathbb{E}_{\mathcal{D}_{n}\sim P_{0}^{\otimes n}}\bigl[\pi\bigl(\{P\in\mathcal{P}:d_{\mathrm{H}}(P,P_{0})\geq M\epsilon_{n,\beta}\}\,\big|\,\mathcal{D}_{n}\bigr)\bigr]$
$\hskip-6.63947pt\overset{(v)}{\leq}C_{1}\epsilon_{n,\beta}^{2}+C_{2}e^{-c_{3}n\epsilon_{n,\beta}^{2}}.$
@@ F.2 Proof of Proposition 6
$g^{\circ}(x)=h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr),$ (57)
@@
Throughout the proof, $C_{1},C_{2},\ldots$ will denote positive quantities that may depend only on $\beta,d,r,c_{0},C_{0}$ and $S$, while $C_{1}^{\prime},C_{2}^{\prime},\ldots$ may in addition depend on $\sigma$. Let $\widetilde{g}_{\beta,r}(x)=\widetilde{g}_{\beta}^{(r)}\bigl(\frac{(U^{(r)})^{\top}x+1_{r}}{2}\bigr)$ be defined as in (10) so that $\widetilde{g}_{\beta,r}\sim\widetilde{\pi}_{\beta,r}$. Observe that $B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})\subseteq B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$ (e.g. Giné and Nickl, 2021, Proposition 4.3.10(ii)). When $\beta<1$, the space $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Hölder space $H^{\beta}$ with equivalent norms (e.g. Giné and Nickl, 2021, Proposition 4.3.23), so for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{1}\|x-y\|_{2}^{\beta}$; when $\beta\geq 1$, $B_{\infty,\infty}^{\beta\wedge 1}$ is equal to the Zygmund space of order 11 with equivalent norms (Triebel, 1983, p. 113), so by Anderson and Pitt (1989, Proposition 2.4) (see also Zygmund, 2002, Chapter II, Theorem 3.4), for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta\wedge 1}([0,1]^{r},2C_{0})$, we have $|g(x)-g(y)|\leq C_{2}\|x-y\|_{2}\bigl\{\log(1/\|x-y\|_{2})\vee 1\bigr\}$. Therefore, for all $x,y\in[0,1]^{r}$ and $g\in B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$,
@@
$\leq\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{(r)})^{\top}x+1_{r}}{2}\biggr)-\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)\biggr|+\biggl|\widetilde{g}_{\beta}^{(r)}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{r}}{2}\biggr)\biggr|$
@@
where the second inequality follows from (61), since $\widetilde{g}_{\beta}^{(r)}\in\mathcal{G}_{\beta}^{(r)}\subseteq B_{\infty,\infty}^{\beta}([0,1]^{r},2C_{0})$ and $\bigl\|\frac{(U^{(r)})^{\top}x+1_{r}}{2}-\frac{(U^{\circ})^{\top}x+1_{r}}{2}\bigr\|_{2}\leq\frac{1}{2}\bigl(n^{-\beta/(2\beta+r)}/\log(en)\bigr)^{1/(\beta\wedge 1)}$ on $E$. Now by Lemma 26, we have